$SPAD/src/input drawex.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

\getchunk{license}
)clear all
-- examples of the draw function
-- graph of a cone
draw(surface(u*cos(v),u*sin(v),u),u = -3..3,v = 0..2*%pi,title == "Cone")
draw(curve(sin(t)*sin(2*t)*sin(3*t),sin(4*t)*sin(5*t)*sin(6*t)),t = 0..2*%pi)
draw(curve(t - sin t,1 - cos t),t = -5..5)
draw(curve(2*t - sin t,2 - cos t),t = -5..5)
draw(curve(t - 2*sin t,1 - 2*cos t),t = -5..5)
draw(surface(5*sin(u)*cos(v),4*sin(u)*sin(v),3*cos(u)),u=0..%pi,v=0..2*%pi)
draw(surface(cos(t)/(1+sin(t)**2),sin(t)*cos(t)*cos(u)/(1+sin(t)**2),
     sin(t)*cos(t)*sin(u)/(1+sin(t)**2)),t = -%pi..%pi,u = 0..%pi)
-- helix
draw(curve(4*cos(t),4*sin(t),t),t = -10..10, title == "Helix")
draw(sin(2 * x**2 + 3 * y**2)/(x**2 + y**2),x = -3..3,y = -3..3)
draw(curve(9*sin(3*t/4),8*sin(t)),t = -4*%pi..4*%pi, _
     title == "Lissajous curve")
draw(curve(-9*sin(4*t/5),8*sin(t)),t = -5*%pi..5*%pi, _
     title == "Lissajous curve")
draw(curve(t**2 + 2*t - 1,t**2 + t - 2),t = -4..3)
draw((x**2 - y**2)/(x**2 + y**2),x = -1..1,y = -1..1)
draw(x**2 - y**2,x = -2..2, y = -2..2)
draw(sin inv x,x = -1.03..3)

draw(sin(x) * sin(y),x = 0..2*%pi, y = 0..2*%pi)
draw(sin(x) * sin(y),x = 20*%pi..22*%pi, y = 20*%pi..22*%pi)
draw(t/100,t = 0..100,coordinates == polar)
draw(cos(x*y),x = -3..3, y = -3..3)
draw(curve(3*(t**2-3),t*(t**2-3)),t = -3..3, title == "Tschirnhausen's Cubic")
draw(curve(sin(t), cos(t), 0), t=0..2*%pi, tubeRadius == 0.5)
draw(curve((t**2-1)/(3*t**2+1),t*(t**2-1)/(3*t**2+1)),t = -3..3, title == _
  "Folium of Descartes")
draw(curve(t - 2*sin t,1 - 2*cos t),t = -5..5)
draw(curve(cos(t)/(1+sin(t)**2),sin(t)*cos(t)/(1+sin(t)**2)),t = _
  -%pi..%pi, title == "Lemniscate of Bernoulli")

———-

References^{[1] nothing}

5